I have done a and b just dont know what formula to use for part c

  1. Norman Osborn, the founder of Oscorp, had a vision of creating a corporation which steadily grows at 5.5% perpetually. Analysts expect Oscorp to pay $4.2 dividend a year from today. The required rate of return from Oscorp shares is 14%.

  1. Calculate the stock price of Oscorp. (2 Marks)
  2. Norman Osborn has been informed by his scientists that they are close to reaching a breakthrough in robotics which will double their annual growth rates in the future. If their dividend grows at the current rate (5.5%) for the next three years (i.e., years 2-4) and the new growth rate is applicable after that, what will be the new stock price of Oscorp? (Note: Oscorp’s next year dividend is $4.2 and the required rate of return is 14%)(5 Marks)

  1. Norman Osborn changed his mind about perpetual dividend payment system. He wants to implement a plan which includes 15 years of dividend payments growing at a certain rate and no dividends afterwards. Furthermore, Norman Osborn does not want the stock price of Oscorp to be affected by this dividend policy change. The Oscorp stocks will pay $5.1 a year from today and the Oscorp stocks are currently priced at 5.5% (perpetual) growth rate. If the required rate of return is still 14%, how much annual growth should Harry aim to successfully implement dividend policy change? (3 Marks)
Answer #1

NOTE: As of you have already solved part (a) and part (b) of the question, the solution provided is for part (c) only

(c) The stock is initially priced by assuming a perpetual growth rate of 5.5 %, expected dividend of $ 5.1 and Required Return of 14%. Norman Osborn wants to replace his dividends in such a way that the stock price remains unaffected.

Unaffected Stock Price = P0 = 5.1 / (0.14-0.055) = $ 60

The new dividends policy intends to pay a constant growth dividend for 15 years and no dividends afterwards.

Let the new constant growth rate be g%

Therefore, 60 = 5.1 x [1/(0.14-g)] x [1-{(1+g)/(1.14)}^(15)]

1.64706 - 11.7647 g = 1 - {(1+g)/(1.14)}^(15)

11.7647 g - 0.64706 = (1+g)^(15) / (1.14)^(15)

83.9757 g - 4.61867 = (1+g)^(15)

Using EXCEL's Goal Seek Function to solve the above equation we get:

g = 0.12151 or 12.151 % 12.15 %